OpenStudy (anonymous):

which equation is equivalent to y - 6 = -12(x + 4)? y = -6x - 48 y = 6x - 48 y = -12x - 42 y = -12x - 54 Please explain! Thanks!

OpenStudy (anonymous):

What do you think the answer is? Expand the brackets and rearrange.

OpenStudy (anonymous):

Honestly, I'm not sure on how to solve these. I missed the class and cant get a recording or anything.

KjSaif (kjsaif):

y−6=−12(x+4) Step 1: Add 6 to both sides. y−6+6=−12x−48+6 y=−12x−42 Answer: y=−12x−42

OpenStudy (anonymous):

No problem, I'll try and explain.\[y-6=-12(x+4)\]Expand the brackets\[y-6=-12*x-12*4\]Add 6 to both sides\[y=-12x-48+6\]\[y=-12x-42\] So the answer is C.

KjSaif (kjsaif):

yeah

OpenStudy (anonymous):

Thanks so much! I've been wondering how to solve these.

KjSaif (kjsaif):

nice job @tom982

KjSaif (kjsaif):

here is a medal